OpenStudy (anonymous):

simplify 75/6 and express as a mixed fraction please show working out as well as answer :D thanks

OpenStudy (anonymous):

Its pretty simple, just take 72 divided by 6 and you get 12 now combine the two as 12 & 2 over 6

OpenStudy (anonymous):

Divide 75 by 6. 6 can divide equally 12 times leaving a remainder of 3. \[\frac{ 75 }{ 6 } = 12\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

thanks guys :D

OpenStudy (anonymous):

@Xpod78 6x12= 72. 75-72= 3. so it will be 3/6 and reduced 1/2

OpenStudy (anonymous):

My bad off by one number :P